# BioDisco: Multi-agent hypothesis generation with dual-mode evidence, iterative feedback and temporal evaluation

### v1

@@ Abstract
Identifying novel hypotheses is essential to scientific research, yet this process risks being overwhelmed by the sheer volume and complexity of available information. Existing automated methods often struggle to generate novel and evidence-grounded hypotheses, lack robust iterative refinement and rarely undergo rigorous temporal evaluation for future discovery potential. To address this, we propose BioDisco, a multi-agent framework that draws upon language model-based reasoning and a dual-mode evidence system (biomedical knowledge graphs and automated literature retrieval) for grounded novelty, integrates an internal scoring and feedback loop for iterative refinement, and validates performance through pioneering temporal and human evaluations and a Bradley–Terry paired comparison model to provide statistically-grounded assessment. Our evaluations demonstrate superior novelty and significance over ablated configurations representative of existing agentic architectures. Designed for flexibility and modularity, BioDisco allows seamless integration of custom language models or knowledge graphs, and can be run with just a few lines of code. We anticipate researchers using this practical tool as a catalyst for the discovery of new hypotheses.
Packagehttps://pypi.org/project/biodisco
Codehttps://github.com/yujingke/BioDisco
@@ 1 Introduction
Knowledge graphs (KGs) provide a structured foundation for storing and analyzing scientific knowledge, with biomedical KGs capturing complex relationships among genes, proteins, diseases, drugs and biological pathways: e.g. HetioNet (Himmelstein et al. 2017) and PharmKG (Zheng et al. 2020). While a variety of tools have been developed to mine such graphs for novel insights (Perdomo-Quinteiro and Belmonte-Hernández 2024), construction and maintenance of KGs are time-consuming and labour-intensive tasks, and a significant proportion of biomedical knowledge remains untapped in unstructured text.
@@
Large language models (LLMs) have emerged as a promising solution, capable of extracting relevant information and identifying latent patterns from vast corpora (Brown et al. 2020), with greater capacity for contextual reasoning than earlier methods of hypothesis generation (c.f. Spangler et al. 2014). However, LLMs risk returning plausible but factually groundless information (Bélisle-Pipon 2024). To mitigate this, models can be imbued with external interfaces, enabling dynamic literature retrieval via retrieval-augmented generation (RAG; Lewis et al. 2020; Singhal et al. 2023).
@@
Evaluating the output of these generative systems—particularly for novel hypothesis generation, in absence of ‘ground truth’—presents its own significant challenge. Whilst general evaluation strategies exist, including expert-based assessments, automated semantic similarity scoring and validation against curated benchmarks (Liu et al. 2025), many recent approaches rely heavily on model-based evaluation, i.e. using other LLMs as judges (Li et al. 2024). Though more accessible than consulting human experts, model-based evaluation is sensitive to the relative strengths of the generator and the judge (Dorner et al. 2024) and alone may not fully capture a system’s capacity for genuine scientific discovery.
Related work
Multi-agent systems
@@
Recent research has explored the use of ‘specialized agents’, often within multi-agent systems, to model the iterative and collaborative nature of scientific discovery. These approaches typically leverage LLMs augmented with tools, assigning distinct roles to individual agents to simulate the research process (Ren et al. 2025). Examples include AI Scientist (Lu et al. 2024), which automates the full research pipeline, ResearchAgent (Baek et al. 2025), a multi-agent system using KGs and peer review-style feedback, and Google’s Co-scientist (Gottweis et al. 2025), which employs a generate–debate–evolve loop. In the biomedical domain, AI agents like Biomni (Huang et al. 2025b) offer generalist capabilities from literature curation to experimental design. More narrowly focussed systems include SciAgents (Ghafarollahi and Buehler 2024) for biomaterials discovery and IntelliScope (Aamer et al. 2025) for biomedical hypothesis generation by identifying meaningful paths in KGs.
Evaluation metrics
@@
Evaluating the efficacy of automated hypothesis generation systems, particularly at scale, involves assessing not only coherence but also novelty, experimental feasibility and potential impact. A prevalent method involves testing a system’s ability to rediscover known scientific findings using historical data (Sybrandt et al. 2018, 2020), though this ‘temporal evaluation’ on held-out data remains underutilized. Assessment by human domain experts is desirable but challenging to perform reliably for large-scale evaluations (Alkan et al. 2025), thus Qi et al. (2024) used LLMs as judges, rating hypotheses for novelty, verifiability, significance, and relevance. This and other LLM evaluations rely on assigning numerical ratings to individual outputs. However, Liusie et al. (2023) showed that paired comparisons using LLMs can outperform direct scoring approaches, at the risk of introducing additional sources of bias, such as order effects. Some authors, e.g. Gottweis et al. (2025), used Elo ratings for automated evaluation. However, a more principled statistical approach would use a probabilistic model to provide uncertainty estimates for ratings and enabling model diagnostics; the Elo system (1978) provides neither.
Benchmarks
@@
Recent efforts further underscore the need for standardized evaluation frameworks. Benchmarks such as PubMedQA (Jin et al. 2019), GPQA (Rein et al. 2024), and CARDBiomedBench (Bianchi et al. 2025) provide question–answer-based datasets to assess LLM reasoning, with a focus on the biomedical domain. Similarly, TruthHypo (Xiong et al. 2025) employs a temporal split to benchmark the classification of associations among biological entities. Complementary initiatives, including HypoBench (Liu et al. 2025), DiscoveryBench (Majumder et al. 2024) and LAB-Bench (Laurent et al. 2024), reinforce the growing emphasis on quantitative assessment of automated hypothesis generation systems. However, most of these benchmarks primarily assess knowledge retrieval, classification, or data-driven insights, falling short of evaluating the generation of novel, multi-entity textual hypotheses that our system aims to produce.
Contributions
@@ 2 Multi-Agent Hypothesis Generator
The BioDisco framework, illustrated in Figure 1, operates through a sophisticated multi-agent architecture, where each agent has a distinct role to facilitate generation, evaluation, refinement and final selection of hypotheses. BioDisco is, to our knowledge, the first system to integrate KG querying, literature search, ‘reviewer’ agents and explicit numerical scoring within a unified agentic feedback loop, wherein each generated hypothesis is internally evaluated against multiple metrics and refined to enhance its quality.
The core of BioDisco’s functionality resides in its network of specialized agents, illustrated in Figure 2. The Background agent searches academic literature to gather scientific evidence, while the Explorer queries a biomedical KG to retrieve relevant subgraphs. The Scientist then formulates initial hypotheses by integrating information from the summarized literature and subgraph data, proposing novel associations. These candidate hypotheses are systematically evaluated by a Critic, who scores each one for novelty, verifiability, relevance and significance (following Qi et al. 2024), providing structured feedback. The Reviewer agent uses this feedback to formulate targeted refinement strategies, such as suggesting new entities or expanded literature review, and invokes dedicated interfaces to conduct more specialized KG queries or fine-grained literature searches. The Refiner amends the hypotheses based on this guidance and additional information. The decision module monitors Critic scores to select and collect hypotheses throughout the feedback loop.
Dual-mode evidence grounding
@@
The KG interface enables dynamic retrieval of structured biomedical knowledge from an external graph. We adopt PrimeKG (Chandak et al. 2023) as the underlying KG, hosted in a Neo4j instance: queries are formulated using Cypher for efficient subgraph retrieval. For each query, agents supply a background summary and a set of keywords. These keywords are first mapped to canonical KG nodes via pretrained biomedical embeddings (BioSimCSEBioLinkBERT-BASE; Kanakarajan et al. 2022), then further filtered for contextual relevance. The resulting set of nodes defines a personalized query, which is executed to retrieve relevant evidence in the form of both direct and multi-hop relations. Query parameters, including relation types and traversal depth, are adaptively set based on task objectives and intermediate feedback. Returned subgraphs are summarized and converted into standardized plaintext representations for downstream use. This approach optimizes retrieval for both efficiency and relevance, enabling flexible and context-aware access to large-scale structured knowledge.
@@
Iterative refinement & multi-agent collaboration
@@ 3 Evaluation of LLM-Generated Hypotheses
This is followed by a model-based ablation study to quantify the contributions of the agentic framework, tools and refinement protocol. Significantly, our evaluation incorporates a pairwise LLM evaluator combined with a Bradley–Terry model. This statistical approach provides more stable and human-aligned assessments of hypothesis quality across multiple dimensions (Liusie et al. 2024), delivering not just point estimates but also uncertainty quantification via 95% comparison intervals, a significant advancement over methods that merely offer direct scores or unidimensional summaries.
@@
Can BioDisco predict future hypotheses?
@@
Datasets
We used two publicly available datasets. (1) An unseen dataset from Qi et al. (2024), containing background–hypothesis pairs curated from biomedical literature, published in August 2023. We filtered this dataset to focus on core biomedical findings, removing entries with psychology or epidemiology backgrounds, resulting in 134 samples. (2) TruthHypo (Xiong et al. 2025) provides keyword pairs across three categories: chemical–gene, gene–gene, and gene–disease. Each pair is labelled with a relationship type: positive, negative, or no relation and was extracted from literature published after 2024. We focused exclusively on positive and negative relations, excluding the ‘no relation’ category, as BioDisco’s primary focus is generation of plausible relationships between entities rather than confirming their absence. From this dataset, we created two subsets. We first selected 30 samples, balanced across categories, to configure a classifier agent. From the remainder, we sampled a test set of 300 instances (limited for computational feasibility), stratified across all categories and the two classes, for evaluation.
Setup
@@
Results
We evaluated the semantic alignment between the generated and ‘gold-standard’ hypotheses from Qi et al. (2024) by computing their cosine similarity using a pretrained biomedical sentence encoder, BioBERTmnli-snli-scinli-scitail-mednli-stsb (Deka et al. 2022). As a baseline, we calculated pairwise similarities between unrelated ‘gold’ hypotheses (i.e. those associated with different background contexts), serving as a negative control for expected similarity in semantically unrelated pairs. Figure 3 shows that generated hypotheses were more semantically similar to the gold standard (median similarity: 0.68) than unrelated gold hypotheses were to each other (median: 0.34), demonstrating that BioDisco reliably produces hypotheses closer to human-curated ground truth than expected by chance.
Category Precision Recall $F_{1}$ Acc.
@@ Results
What are the effects of knowledge & refinement?
@@
Leveraging the established correlation between LLM-based evaluators and human expert judgements (Lu et al. 2024), we employed a dedicated pairwise LLM evaluator, rather than direct scoring of individual hypotheses.
Each of the systems was given 100 inputs spanning a diversity of topics, from oncology to ageing and cellular senescence, generating one hypothesis per input. For every pair of generated hypotheses from different configurations, an LLM judge selected its preferred candidate (or declared a tie) for each of four metrics: novelty, relevance, verifiability and significance (from Qi et al. 2024).
@@
$\log\operatorname{odds}(i~\text{beats}~j)=\alpha+\beta_{i}-\beta_{j},$ (1)
where $i$ and $j$ denote ‘players’ or LLM configurations, $\beta_{i}$ is the ability score of the $i$th LLM system and $\alpha$ is a ‘home advantage’ parameter, estimating the possible order effect bias, assuming the hypothesis of $i$ is presented to the evaluator first in the pair. The continuous latent ability scores for each configuration and each metric provide a univariate ranking that captures relative performance.
Figure 4: Centipede plot of ability scores for BioDisco and four ablation configurations, with 95% comparison intervals. A multi-agent system clearly outperforms a single LLM (GPT-4.1) generating novel, significant hypotheses; tool use (i.e. KG and literature search) and iterative refinement each yield further improvements
The fitted ability scores and their 95% comparison intervals (calculated using a quasi-variance approximation; Firth 2004) are visualized in Figure 4. The model shows a clear, progressive improvement in hypothesis novelty and significance as components are added to the system. The single LLM baseline was least preferred. Introducing a multi-agent structure provided significant improvement, suggesting a benefit from role-based reasoning alone. Performance further increased with the addition of external tools and with iterative refinement. Between the two, tool use provided a greater benefit, emphasizing the value of grounding the generation process in external knowledge sources. Finally, the full BioDisco system was most preferred for these metrics, demonstrating a synergistic effect between dual-mode evidence and iterative refinement, validating our approach.
@@
Additional results—including order effect estimates and a direct scoring evaluation—are provided in the Appendix.
What do human domain experts think?

### v2

@@ Abstract
Identifying novel hypotheses is essential to scientific research, yet this process risks being overwhelmed by the sheer volume and complexity of available information. Existing automated methods often struggle to generate novel and evidence-grounded hypotheses, lack robust iterative refinement and rarely undergo rigorous temporal evaluation for future discovery potential. To address this, we propose BioDisco, a multi-agent framework that draws upon language model-based reasoning and a dual-mode evidence system (biomedical knowledge graphs and automated literature retrieval) for grounded novelty, integrates an internal scoring and feedback loop for iterative refinement, and validates performance through pioneering temporal and human evaluations and a Bradley–Terry paired comparison model to provide statistically-grounded assessment. Our evaluations demonstrate superior novelty and significance over ablated configurations and generalist biomedical agents. Designed for flexibility and modularity, BioDisco allows seamless integration of custom language models or knowledge graphs, and can be run with just a few lines of code.
@@ 1.1 Background
Knowledge graphs (KGs) provide a structured foundation for storing and analyzing scientific knowledge, with biomedical KGs capturing complex relationships among genes, proteins, diseases, drugs and biological pathways: e.g. HetioNet (Himmelstein et al., 2017) and PharmKG (Zheng et al., 2020). While a variety of tools have been developed to mine such graphs for novel insights (Perdomo-Quinteiro and Belmonte-Hernández, 2024), construction and maintenance of KGs are time-consuming and labour-intensive tasks, and a significant proportion of biomedical knowledge remains untapped in unstructured text.
@@
Large language models (LLMs) have emerged as a promising solution, capable of extracting relevant information and identifying latent patterns from vast corpora (Brown et al., 2020), with greater capacity for contextual reasoning than earlier methods of hypothesis generation (c.f. Spangler et al., 2014). However, LLMs risk returning plausible but factually groundless information (Bélisle-Pipon, 2024). To mitigate this, models can be imbued with external interfaces, enabling dynamic literature retrieval via retrieval-augmented generation (RAG; Lewis et al., 2020; Singhal et al., 2023).
@@
Evaluating the output of these generative systems—particularly for novel hypothesis generation, in absence of ‘ground truth’—presents its own significant challenge. Whilst general evaluation strategies exist, including expert-based assessments, automated semantic similarity scoring and validation against curated benchmarks (Liu et al., 2025), many recent approaches rely heavily on model-based evaluation, i.e. using other LLMs as judges (Li et al., 2024). Though more accessible than consulting human experts, model-based evaluation is sensitive to the relative strengths of the generator and the judge (Dorner et al., 2024) and alone may not fully capture a system’s capacity for genuine scientific discovery.
1.2 Related work
1.2.1 Multi-agent systems
@@
Recent research has explored the use of ‘specialized agents’, often within multi-agent systems, to model the iterative and collaborative nature of scientific discovery. These approaches typically leverage LLMs augmented with tools, assigning distinct roles to individual agents to simulate the research process (Ren et al., 2025). Examples include AI Scientist (Lu et al., 2024), which automates the full research pipeline, ResearchAgent (Baek et al., 2025), a multi-agent system using KGs and peer review-style feedback, and Google’s Co-scientist (Gottweis et al., 2025), which employs a generate–debate–evolve loop. In the biomedical domain, AI agents like Biomni (Huang et al., 2025b) offer generalist capabilities from literature curation to experimental design. More narrowly focussed systems include SciAgents (Ghafarollahi and Buehler, 2024) for biomaterials discovery and IntelliScope (Aamer et al., 2025) for biomedical hypothesis generation by identifying meaningful paths in KGs.
1.2.2 Evaluation metrics
@@
Evaluating the efficacy of automated hypothesis generation systems, particularly at scale, involves assessing not only coherence but also novelty, experimental feasibility and potential impact. A prevalent method involves testing a system’s ability to rediscover known scientific findings using historical data (Sybrandt et al., 2018; 2020), though this ‘temporal evaluation’ on held-out data remains underutilized. Assessment by human domain experts is desirable but challenging to perform reliably for large-scale evaluations (Alkan et al., 2025), thus Qi et al. (2024) used LLMs as judges, rating hypotheses for novelty, verifiability, significance, and relevance. This and other LLM evaluations rely on assigning numerical ratings to individual outputs. However, Liusie et al. (2023) showed that paired comparisons using LLMs can outperform direct scoring approaches, at the risk of introducing additional sources of bias, such as order effects. Some authors, e.g. Gottweis et al. (2025), used Elo ratings for automated evaluation. However, a more principled statistical approach would use a probabilistic model to provide uncertainty estimates for ratings and enabling model diagnostics; the Elo system (1978) provides neither.
1.2.3 Benchmarks
@@
Recent efforts further underscore the need for standardized evaluation frameworks. Benchmarks such as PubMedQA (Jin et al., 2019), GPQA (Rein et al., 2024), and CARDBiomedBench (Bianchi et al., 2025) provide question–answer-based datasets to assess LLM reasoning, with a focus on the biomedical domain. Similarly, TruthHypo (Xiong et al., 2025) employs a temporal split to benchmark the classification of associations among biological entities. Complementary initiatives, including HypoBench (Liu et al., 2025), DiscoveryBench (Majumder et al., 2024) and LAB-Bench (Laurent et al., 2024), reinforce the growing emphasis on quantitative assessment of automated hypothesis generation systems. However, most of these benchmarks primarily assess knowledge retrieval, classification, or data-driven insights, falling short of evaluating the generation of novel, multi-entity textual hypotheses that our system aims to produce.
1.3 Contributions
@@ 2 Multi-Agent Hypothesis Generator
The BioDisco framework, illustrated in Figure 1, operates through a sophisticated multi-agent architecture, where each agent has a distinct role to facilitate generation, evaluation, refinement and final selection of hypotheses. BioDisco is, to our knowledge, the first system to integrate KG querying, literature search, ‘reviewer’ agents and explicit numerical scoring within a unified agentic feedback loop, wherein each generated hypothesis is internally evaluated against multiple metrics and refined to enhance its quality. A comparison with other systems is presented in Table 1.
The core of BioDisco’s functionality resides in its network of specialized agents, illustrated in Figure 2. The Background agent searches academic literature to gather scientific evidence, while the Explorer queries a biomedical KG to retrieve relevant subgraphs. The Scientist then formulates initial hypotheses by integrating information from the summarized literature and subgraph data, proposing novel associations. These candidate hypotheses are systematically evaluated by a Critic, who scores each one for novelty, verifiability, relevance and significance (following Qi et al., 2024), providing structured feedback. The Reviewer agent uses this feedback to formulate targeted refinement strategies, such as suggesting new entities or expanded literature review, and invokes dedicated interfaces to conduct more specialized KG queries or fine-grained literature searches. The Refiner amends the hypotheses based on this guidance and additional information. The decision module monitors Critic scores to select and collect hypotheses throughout the feedback loop.
Table 1: Features of LLM agentic systems: multiple agents, use of external tools, knowledge graph integration, ability (dynamically) to search academic literature, presence of a ‘reviewer’ agent critically appraising outputs, and refinement of hypotheses based on explicit numerical scoring. ResearchAgent (Baek et al., 2025) constructs a knowledge graph from an academic graph, but does not access a ‘live’ search API
Co-scientist Intelliscope SciAgents ResearchAgent BioDisco
Multi-agent
Tool use
Knowledge graph
Search
Reviewer
Scoring
2.1 Dual-mode evidence grounding
@@
The KG interface enables dynamic retrieval of structured biomedical knowledge from an external graph. We adopt PrimeKG (Chandak et al., 2023) as the underlying KG, hosted in a Neo4j instance: queries are formulated using Cypher for efficient subgraph retrieval. For each query, agents supply a background summary and a set of keywords. These keywords are first mapped to canonical KG nodes via pretrained biomedical embeddings (BioSimCSEBioLinkBERT-BASE; Kanakarajan et al., 2022), then further filtered for contextual relevance. The resulting set of nodes defines a personalized query, which is executed to retrieve relevant evidence in the form of both direct and multi-hop relations. Query parameters, including relation types and traversal depth, are adaptively set based on task objectives and intermediate feedback. Returned subgraphs are summarized and converted into standardized plaintext representations for downstream use. This approach optimizes retrieval for both efficiency and relevance, enabling flexible and context-aware access to large-scale structured knowledge.
@@
2.2 Iterative refinement & multi-agent collaboration
@@ 3 Evaluation of LLM-Generated Hypotheses
This is followed by a model-based ablation study to quantify the contributions of the agentic framework, tools and refinement protocol. Significantly, our evaluation incorporates a pairwise LLM evaluator combined with a Bradley–Terry model. This statistical approach provides more stable and human-aligned assessments of hypothesis quality across multiple dimensions (Liusie et al., 2024), delivering not just point estimates but also uncertainty quantification via 95% comparison intervals, a significant advancement over methods that merely offer direct scores or unidimensional summaries.
@@
3.1 Can BioDisco predict future hypotheses?
@@
3.1.1 Datasets
We used two publicly available datasets. (1) An unseen dataset from Qi et al. (2024), containing background–hypothesis pairs curated from biomedical literature, published in August 2023. We filtered this dataset to focus on core biomedical findings, removing entries with psychology or epidemiology backgrounds, resulting in 134 samples. (2) TruthHypo (Xiong et al., 2025) provides keyword pairs across three categories: chemical–gene, gene–gene, and gene–disease. Each pair is labelled with a relationship type: positive, negative, or no relation and was extracted from literature published after 2024. We focused exclusively on positive and negative relations, excluding the ‘no relation’ category, as BioDisco’s primary focus is generation of plausible relationships between entities rather than confirming their absence. From this dataset, we created two subsets. We first selected 30 samples, balanced across categories, to configure a classifier agent. From the remainder, we sampled a test set of 300 instances (limited for computational feasibility), stratified across all categories and the two classes, for evaluation.
3.1.2 Setup
@@
3.1.3 Results
We evaluated the semantic alignment between the generated and ‘gold-standard’ hypotheses from Qi et al. (2024) by computing their cosine similarity using a pretrained biomedical sentence encoder, BioBERTmnli-snli-scinli-scitail-mednli-stsb (Deka et al., 2022). As a baseline, we calculated pairwise similarities between unrelated ‘gold’ hypotheses (i.e. those associated with different background contexts), serving as a negative control for expected similarity in semantically unrelated pairs. Figure 3 shows that generated hypotheses were more semantically similar to the gold standard (median similarity: 0.68) than unrelated gold hypotheses were to each other (median: 0.34), demonstrating that BioDisco reliably produces hypotheses closer to human-curated ground truth than expected by chance.
Category Precision Recall $F_{1}$ Accuracy
@@ 3.1.3 Results
3.2 Ablation of knowledge access & refinement
@@
For a direct comparison with other biomedical agentic systems, we also generated a series of hypotheses with Biomni (Huang et al., 2025b):
  1. 6.
    Biomni: a general-purpose biomedical AI agent, based on GPT-4.1, with access to various bioinformatics tools but no predefined workflow or iteration loop.
Leveraging the established correlation between LLM-based evaluators and human expert judgements (Lu et al., 2024), we employed a dedicated pairwise LLM evaluator, rather than direct scoring of individual hypotheses.
Each of the systems was given 100 inputs spanning a diversity of topics, from oncology to ageing and cellular senescence, generating one hypothesis per input. For every pair of generated hypotheses from different configurations, an LLM judge selected its preferred candidate (or declared a tie) for each of four metrics: novelty, relevance, verifiability and significance (from Qi et al., 2024).
@@
$\log\operatorname{odds}(i~\text{beats}~j)=\beta_{i}-\beta_{j}+\alpha z_{ij},$ (1)
where $i$ and $j$ denote ‘players’ or LLM configurations, $\beta_{i}$ is the ability score of the $i$th LLM system and $\alpha$ is a ‘home advantage’ parameter, estimating the possible order effect bias, with $z_{ij}=1$ if the hypothesis of $i$ is presented to the evaluator first in the pair, and $z_{ij}=-1$ otherwise. Ties are treated as a half-win for each player. The continuous latent ability scores for each configuration and each metric provide a univariate ranking that captures relative performance.
Figure 4: Centipede plot of ability scores for BioDisco, four ablation configurations and Biomni, with 95% comparison intervals. A multi-agent system clearly outperforms a single LLM (GPT-4.1) generating novel, significant hypotheses; tool use (i.e. KG and literature search) and iterative refinement each yield further improvements. Biomni, a competing system, is mostly better than a single LLM and produces verifiable hypotheses, but less novel and significant than the full BioDisco framework
The fitted ability scores and their 95% comparison intervals (calculated using a quasi-variance approximation; Firth, 2004) are visualized in Figure 4. The model shows a clear, progressive improvement in hypothesis novelty and significance as components are added to the system. The single LLM baseline was least preferred. Introducing a multi-agent structure provided significant improvement, suggesting a benefit from role-based reasoning alone. Performance further increased with the addition of external tools and with iterative refinement. Between the two, tool use provided a greater benefit, emphasizing the value of grounding the generation process in external knowledge sources. Finally, the full BioDisco system was most preferred for these metrics, demonstrating a synergistic effect between dual-mode evidence and iterative refinement, validating our approach.
@@
Estimates for the order effect parameter $\hat{\alpha}$ are given in Table 3 and suggest a statistically significant order bias for Verifiability evaluations (in favour of the item appearing first in the comparisons) but not enough evidence to indicate such an effect exists in Novelty, Relevance or Significance comparisons at the 5% level of significance.
Metric Estimate Std. Error Statistic $p$-value
Novelty 0.10-0.10 0.21 0.51-0.51 0.61
Relevance 0.26 0.11 2.38 0.02
Significance 0.09 0.17 0.52 0.60
Verifiability 1.11 0.13 8.68 0.00
Additional results—including a direct scoring evaluation—are provided in the Appendix.
3.3 Human expert evaluation